OpenStudy (anonymous):

Question 2: Write the equation of the line in slope-intercept form given the point (4,3) and the slope of 1/2.

hartnn (hartnn):

since the point and slope is given, we use slope point form first, the point slope form of equation of line is \(\large y-y_1=m(x-x_1)\) where m=slope =1/2 here. point =(x1,y1)= (4,3) --->x1=4 ,y1=3 just plug in values, to get the equation of line first, then i'll tell you how to convert it into slope intercept form